For the next three questions, consider a closed economy in which the money supply totals $9000, growing at a rate of 2.2%. Velocity is constant at a value of 4, and the price level has a growth rate of 1%, currently at a value of 1 as well.  

1) Carefully following all instructions, calculate real output for this economy.

2) Carefully following all instructions, calculate the growth rate of real output for this economy.

3) Now let's say everything else remains the same, but now the economy has real output of 54,000 units. Let's further say only velocity can adjust. What will the new value of velocity be?
Answer #1

Money supply Ms = 9000. Money supply growth rate = 2.2% Velocity = 4 and growth rate of velocity = 0%. Inflation rate = 1%. Price level = 1.

1) Using quantity theory we have P * Y = Ms * V

1 * Y = 9000 * 4

Hence real output is 36000

2) Growth rate of output = growth rate of money supply + growth rate of velocity of money - rate of inflation

= 2.2% + 0% - 1%

= 1.2%

3) Now we have P x Y = Ms x V

1 * 54000 = 9000 * V

V = 54000/9000 = 6

Hence velocity rises to 6.